Can you please help me with this problem?
How many gallons of a 10% acid solution must be mixed with 5 gal of 25% acid solution to make a 15% acid solution?

Answer by mananth(16946)

-------------- percent ---------------- Amount
Acid solution I 25 ---------------- 5 gallons
Acid solution II 10 ---------------- x gallons
Total 15 ---------------- 5 + x gallons

25 * 5 + 10 x = 15 ( 5 + x )
125 + 10 x = 75 + 15 x
10 x -15 x = -125 + 75
-5 x = -50
/ -5
x = 10 of gallons Acid solution II
10 of gallons Acid solution I
m.ananth@hotmail.ca